OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)] −13 −25 25 17

OpenStudy (anonymous):

(f.g)(x) = -2x² + x + 6 ⇒ (f.g)(6) = -2 * 36 + 6 + 6 = -72 + 12 = -60

OpenStudy (anonymous):

?

OpenStudy (anonymous):

thats not even one of the answers

OpenStudy (anonymous):

lol i know hold on one sec and i will do it

OpenStudy (anonymous):

g(-2) = 6 * (-2) - 1 = -12 - 1 = -13 ⇒ f(g(-2)) = f(-13) = -2 * (-13) - 1 = 26 - 1 = 25 

OpenStudy (anonymous):

their is you answer

OpenStudy (anonymous):

Ty that what I thought it was.

OpenStudy (anonymous):

ok thank you have a good day